I see similar questions, but not quite the same and am having problems:
How many ounces of a 30% gold solution must be mixed with 60 ounces of a 15% solution to obtain a solution that is 20% gold?
The teacher released the answer as 30oz. of the 10% solution....not sure how to get there

Answer by ptaylor(1894)


Let x=amount of 30% solution needed
Amount of pure gold before the mixture(0.30x+0.15*60) has to equal the amount of pure gold after the mixture (0.20(60+x)), so
0.30x+0.15*60=0.20(60+x) get rid of parens and simplify
0.30x+9=12+0.20x subtract 9 and also 0.20x feom each side
0.30x-0.20x+9-9=0.20x-0.20x+12-9
0.10x=3 divide each side by 0.10
x=30 oz-------------------amount of 30% solution needed
CK
0.30*30+0.15*60=0.20*90
9+9=18
18=18
Hope this helps---ptaylor